These were accepted.
Alastair reported that there seemed to be some issues regarding the funding of the research data storage project (CCPAG report). There may be knock-on effects on EdLAN 10Gbps connectivity options.
This action transferred to Chris, as documentation editor.
There are now only two "220-kernel" machines remaining.
... but ALL still to populate the page
Toby has incorporated suggestions. Alastair will take the revised version to CSG.
Toby has a revised list
Report from Computing Executive Group
Reports from units
Yet another possibility would be to warn in advance then do option "2" during a lunch-time or similarly-quiet period. Inf-unit will give it some more thought.
Schedule some tests.
Toby will be following this up on the Heimdal list. It would be nice to resolve it, but it's not worth putting a lot of effort into it.
Iain raised the issue of School-bought laptops coming with 32-bit Windows 7, as 64-bit seemed not to be an option, at least for the model he had. Alastair will check on this and discuss with Sheila.
Graham to discuss the issues with MP-unit.
There might not be a <stable> release during the week 25th-29th March unless something crops up. May 1st's might be moved to Thursday.
This prompted some discussion. The overall conclusion was that we should check with the Institute before trying to fix out-of-warranty machines. There would be some merit in warning Institutes when their machines were about to go OOW, assuming the right people could be found, though some doubt as to how effective it would be in general. We should encourage Institutes not to keep valuable data on machines which might be unreasonably hard to fix or have risky storage configurations.
This is the last remaining VMware machine. New tickets are going into the forum-issues queue on rt4.inf. Old tickets are proving hard to move. If it isn't possible to find a way to move them in the next couple of days they'll be put into a RT3 running on SL6 until such time as the remaining live ones are resolved.
Group 1: Craig, Ian, Iain, Jennifer, Richard has worked one day so far.
Group 2: Neil, Alastair, Toby, Ross has finished.
Group 3: Roger, Chris, George, Graham, Lindsey has finished.
Group 4: Gordon, Stephen, Tim, Alison, Carol has worked one day so far.
All groups should be done by next month's Development meeting.
Topics for discussion
Proposal to Restructure the LCFG headers (briefing paper). This prompted much discussion:
The next meeting will be on Wednesday 27th March at 10am in IF-4.31, chaired by Craig.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh